# Climate-Invariant Machine Learning

### 06 Jan 2024 v4

@@ 1.1 Background
In climate applications, this extrapolation issue means that ML algorithms typically fail when exposed to dynamic, thermodynamic, or radiative conditions that differ substantially from the range of conditions they were trained on. Examples include [27], who showed that an RF-based moist convection scheme generalizes poorly in the tropics of a climate 6.5K warmer than the training climate; and [28], who showed that NNs and support vector machines downscaling surface air temperature made significant extrapolation errors when exposed to temperatures 2-3K warmer than in the training set. [29] showed that an NN-based thermodynamic subgrid-scale closure generalizes well to climates 1-2K warmer than the training one, but makes large errors as soon as the test climate is 4K warmer than the training one. [30] confirmed that these generalization errors remain even when the NN subgrid closure is modified to enforce conservation laws to within machine precision. This has led several studies to recommend training ML models in multiple climates if possible [31, 32]. Interestingly, both Guillaumin et al. [33] who trained an NN parameterization for subgrid oceanic momentum transport and Molina et al. [34] who trained convolutional NNs [35] to classify thunderstorms in high-resolution model outputs found that their ML models generalized well to a warmer climate. While this may be because both models relied heavily on velocity inputs and their gradients, whose distributions changed only slightly when the climate warmed, Molina et al. noted that using two types of ML layers, namely batch normalization (BN, [36]) followed by dropout (DP, [37]), was key to this successful generalization.
@@ 2 Data
with the exception of SAM for which only the (-4K) and (+0K) simulations are available. By prescribing surface temperature, we focus on ML’s ability to consistently predict the atmospheric response to climate change across configurations. Projecting climate change involves a broader range of processes and is beyond this work’s scope. We summarize the simulations and indicate their spatiotemporal resolutions in Tab S1. Fig 2 gives a visualization of surface temperatures in each model (Fig 2a) and snapshots of mid-tropospheric subgrid heating (Fig 2b), which is one of our ML models’ outputs.
@@ 2.1 Super-Parameterized Aquaplanet Simulations
We use data from two-year SPCAM3 [56] climate simulations in an aquaplanet configuration [57], with zonally-symmetric surface temperatures fixed to a realistic meridionally-asymmetric profile [58]. The insolation is fixed to boreal summer conditions with a full diurnal cycle. A two-dimensional storm-resolving model is embedded in each grid cell of SPCAM3, namely 8 SAM atmospheric columns using a spatiotemporal resolution of 4km30lev25s, and the default one-moment microphysical scheme [59]. SPCAM3 combines a spectral primitive equation solver with a semi-Lagrangian dynamical core for advection [57]. The (+0K) SPCAM3 simulation was first presented in [53] and subsequently used to train ML subgrid closures in [29, 50, 60]. Inspired by the generalization experiment of [27], the (+4K) simulation was introduced in [29], and we ran the (-4K) simulation for the work presented here to increase the surface temperature generalization gap from 4K to 8K.
@@ 2.4 Normalization and Training, Validation, and Test Split
Both generalization experiments expose ML models to out-of-distribution inputs they have not been trained on. Following best ML practices [70], we use the training set to optimize the ML model’s trainable parameters, save the trainable parameters that led to the best performance on the validation set to avoid overfitting the training set, and evaluate the final model on samples from a separate test set. We split each of the 9 simulations into training/validation/test sets by using non-contiguous 3-month periods (reported in Tab S1) to avoid high temporal correlations between training/validation/test set samples [71]. Following [29], the normalization procedure involves subtracting the mean value of each input variable at each vertical level and dividing by the maximum range of that variable across the entire atmospheric column.
@@ 3 Theory
Throughout the following section, we compare two transformation options for each input, whose univariate PDFs are depicted for all three atmospheric models in Fig 3: No transformation (top) and our most successful transformation (bottom). All transformations are derived in SM B2. Our comparison relies on the Hellinger and Jensen-Shannon PDF distance metrics defined and calculated in Sec 6. To prevent information leaks from generalization test sets into the physically-informed ML framework, we take two precautions: (1) the physical transformations are fixed, meaning that their structure and parameters are non-trainable; and (2) transformations are ranked based on their generalization from (-4K) to (+0K) in SPCAM3. Our (+4K) results across models and configurations independently confirm this ranking.

### 17 Jan 2024 v5

@@ 1.1 Background
In climate applications, this extrapolation issue means that ML algorithms typically fail when exposed to dynamic, thermodynamic, or radiative conditions that differ substantially from the range of conditions they were trained on. Examples include [27], who showed that an RF-based moist convection scheme generalizes poorly in the tropics of a climate 6.5K warmer than the training climate; and [28], who showed that NNs and support vector machines downscaling surface air temperature made substantial extrapolation errors when exposed to temperatures 2-3K warmer than in the training set. [29] showed that an NN-based thermodynamic subgrid-scale closure generalizes well to climates 1-2K warmer than the training one, but makes large errors as soon as the test climate is 4K warmer than the training one. [30] confirmed that these generalization errors remain even when the NN subgrid closure is modified to enforce conservation laws to within machine precision. This has led several studies to recommend training ML models in multiple climates if possible [31, 32]. Interestingly, both Guillaumin et al. [33] who trained an NN parameterization for subgrid oceanic momentum transport and Molina et al. [34] who trained convolutional NNs [35] to classify thunderstorms in high-resolution model outputs found that their ML models generalized well to a warmer climate. While this may be because both models relied heavily on velocity inputs and their gradients, whose distributions changed only slightly when the climate warmed, Molina et al. noted that using two types of ML layers, namely batch normalization (BN, [36]) followed by dropout (DP, [37]), was key to this successful generalization.
@@ 2 Data
with the exception of SAM for which only the (-4K) and (+0K) simulations are available. By prescribing surface temperature, we focus on ML’s ability to consistently predict the atmospheric response to climate change across configurations. Projecting climate change involves a broader range of processes and is beyond this work’s scope. We summarize the simulations and indicate their spatiotemporal resolutions in Tab S1. Fig 2 gives a visualization of surface temperatures in each model and Fig S1 provides snapshots of mid-tropospheric subgrid heating (Fig 2b), which is one of our ML models’ outputs.
@@ 2.1 Super-Parameterized Aquaplanet Simulations
We use data from two-year SPCAM3 [56] climate simulations in an aquaplanet configuration [57], with zonally-symmetric surface temperatures fixed to a realistic meridionally-asymmetric profile [58]. The insolation is fixed to boreal summer conditions with a full diurnal cycle. A two-dimensional storm-resolving model is embedded in each grid cell of SPCAM3, namely 8 SAM atmospheric columns using a spatiotemporal resolution of 4km25s, and the default one-moment microphysical scheme [59]. SPCAM3 combines a spectral primitive equation solver with a semi-Lagrangian dynamical core for advection [57]. The (+0K) SPCAM3 simulation was first presented in [53] and subsequently used to train ML subgrid closures in [29, 50, 60]. Inspired by the generalization experiment of [27], the (+4K) simulation was introduced in [29], and we ran the (-4K) simulation for the work presented here to increase the surface temperature generalization gap from 4K to 8K.
@@ 2.4 Normalization and Training, Validation, and Test Split
Both generalization experiments expose ML models to out-of-distribution inputs they have not been trained on. Following best ML practices [70], we use the training set to optimize the ML model’s trainable parameters, save the trainable parameters that led to the best performance on the validation set to avoid overfitting the training set, and evaluate the final model on samples from a separate test set. We split each of the 8 simulations into training/validation/test sets by using non-contiguous 3-month periods (reported in Tab S1) to avoid high temporal correlations between training/validation/test set samples [71]. Following [29], the normalization procedure involves subtracting the mean value of each input variable at each vertical level and dividing by the maximum range of that variable across the entire atmospheric column.
@@ 3 Theory
Throughout the following section, we compare two transformation options for each input, whose univariate PDFs are depicted for all three atmospheric models in Fig 3: No transformation (top) and our most successful transformation (bottom). All transformations are derived in SM B2. Our comparison relies on the Hellinger and Jensen-Shannon PDF distance metrics defined and calculated in Sec 6 and SM D1. To prevent information leaks from generalization test sets into the physically-informed ML framework, we take two precautions: (1) the physical transformations are fixed, meaning that their structure and parameters are non-trainable; and (2) transformations are ranked based on their generalization from (-4K) to (+0K) in SPCAM3. Our (+4K) results across models and configurations independently confirm this ranking.